Category Archives: MALDI-TOF status

Tania

February 12, 2015

I can’t believe it’s been 2 weeks since the last post! We’ve been so busy! Today we are recovering from one of our electrical shutdowns: Orbi is baking out, MALDI is ready and is being used by Krista from the Meredith group.

Here’s the current Orbi queue:  Katsu, Nick, Huihun, Joana, Jung (49 hours). MALDI queue: Alex, Jason, Core Facility standards (3 hours).

If Orbi behaves, all analyses should be done by Tuesday next week.

Tania

January 28, 2015

Thanks to the OPP electricians, we have our power restored, and the instruments are pumping down. Orbi will take some time to reach operating conditions, as usual. MALDI is ready for your samples. LCT and QTRAP should be ready tomorrow.

Meanwhile, Scott’s, Joana’s, and John’s gels are incubating with trypsin overnight.

Tania

January 23, 2015

Orbi is running Olga’s second sample, and Ninad’s two samples are next in the queue. Gaozhong’s, Ming-Yang’s, and Yang’s samples have been analyzed and the data will be processed on Monday.

I have analyzed Alex’s oligos using MALDI TOF, next will be polymers for Brooks on Monday.

Tania

January 16, 2015

The instrument was not performing to the specs during the past 2 weeks, but it looks like the issue has been resolved with the help of two Thermo tech support specialists, Mike Henry and Mohit Jain. I will check the calibrations next week to make sure everything is to the specs.

Meanwhile, Orbi is analyzing Yuning’s digest. Ninad’s samples are next. I have cleaned all gel bands submitted to date, and they are currently incubating with trypsin.

MALDI-TOF is working. I ran polymer samples for Brooks. Next are oligonucleotides for Alex and milk samples for Asha.

Tania

December 22, 2014

Orbi is running second set of Scott’s samples; the first set has been analyzed and the data processed. Next in the queue are Eric’s and Josh’s samples, after which the instrument will be shut down for the holidays.

MALDI will also be shut down tomorrow to avoid potential problems (based on the last year’s holidays experience).

 

Tania

December 10, 2014

Orbi is running Ninad’s IPs. Eric’s samples are next. Scott’s gel lanes will be ready for proteolysis by tomorrow morning. We have heating, cooling, and electricity, all at the same time!

MALDI is experiencing severe deficit of your attention.

Tania

December 8, 2014

We are two weeks away from the winter holiday break (12/22 – 01/04), during which Orbi will be down. This morning, I received 48 samples for Orbi; so counting the 24 samples currently running, plus 6 more waiting, the Orbi queue is completely full for the 2 weeks before the break. I ask you to keep your samples safe in your refrigerators or freezers until we return in January.

MALDI currently doesn’t have a queue.

Tania

December 4, 2014

We had a power interruption today to allow electricians repair a faulty circuit breaker. All mass spectrometers were down, and now we are getting back to operating conditions. This will cause some delays in your sample analyses.

I should be able to resume running Yuanjun’s samples on Monday, after Orbi gets stable temperature so I can calibrate it.

MALDI is ready.

LCT is pumping down.

QTrap is pumping down.